OpenStudy (anonymous):

Solve for value of tan3pi/16 \[1+\sqrt2=\frac{ 2\tan \frac{ 3\pi }{ 16 } }{ 1-\tan^2\frac{ 3\pi }{ 16 } }\]

geerky42 (geerky42):

So basically you just have \(1+\sqrt2=\dfrac{2x}{1-x^2}\)

OpenStudy (anonymous):

So I move the denominator to the left side?

geerky42 (geerky42):

Yeah I would do that. \((1+\sqrt2)(x^2-1) = 2x\)

geerky42 (geerky42):

Arrange a little more, then you can apply quadratic formula, I guess.

OpenStudy (anonymous):

Do I expand it first?

OpenStudy (anonymous):

How can I get the answer to exact values?

geerky42 (geerky42):

Nah just distribute \((1+\sqrt2)\) You want equation to be in form \(ax^2+bx+c=0\) So here \(-(1+\sqrt2)x^2-2x+(1+\sqrt2)\) So now we can apply quadratic formula: \(a=-(1+\sqrt2)\\b=-2\\c=1+\sqrt2\)

geerky42 (geerky42):

\[x = \dfrac{-b+\sqrt{b^2-4ac}}{2a}\]

zepdrix (zepdrix):

Hmm this process seems pretty beastly :) Do you remember the Tangent Double Angle Formula we used in the last problem Sammy? You probably should start with that. Just another idea.

OpenStudy (anonymous):

I'm not getting this.. :(

geerky42 (geerky42):

Hmm, yeah this method is much more preferable. Tangent Double Angle Formula: \(\displaystyle \tan(2x) = \dfrac{2\tan(x)}{1-\tan^2(x)}\)

OpenStudy (anonymous):

What do I do with the tan(2x) identity?

OpenStudy (anonymous):

Oh I got it now! @zepdrix :) I kinda used the quadratic formula. I took me a long time to figure it out!

zepdrix (zepdrix):

Oh you figure it out? :D Sorry I thought pony was helping you.